OpenStudy (anonymous):

help!!!needed 2.24 L methane gas is mixed with 1.12 L of ethane gas in cylinder with movable piston at SATP where pressure is 1 atm and temperature is 298 k.The total number of carbon atoms present in this mixture at SATP is?

OpenStudy (anonymous):

@hartnn @ganeshie8

OpenStudy (masumanwar):

pv =nRT this equation may help to solve this problem

OpenStudy (anonymous):

I tried..can you elaborate it....

OpenStudy (aaronq):

what did you do so far?

OpenStudy (anonymous):

I didn't understand the question.... @aaronq

OpenStudy (aaronq):

you said you tried using the ideal gas law. What you wanna do is find the moles of each gas separately using PV=nRT Then use the moles to find the number of carbon atoms

OpenStudy (anonymous):

explanation please...!! @aaronq

OpenStudy (aaronq):

have you tried doing it yourself?

OpenStudy (anonymous):

yes ....

OpenStudy (anonymous):

@chmvijay

OpenStudy (chmvijay):

can u show me what you did ?

OpenStudy (somy):

Step 1 Use PV=nRT formula to find mole of EACH methane gas and ethane gass

OpenStudy (somy):

n=PV/RT P= 1 atm V1= 2.24 for methane V2= 1.12 for ethane R= constant value 0.08206 T= 298K so now solve this using V1 meaning methane value and get the mole then solve it again but now using V2 meaning ethane value and get ethane mole

OpenStudy (somy):

Step 2 after getting mole of both now our target is mole of C in methane and ethane so knowing mole of methane we can do this mathematically like this : mole of methane = molecular mass of methane Unknown mole of Carbon = molecular mass of Carbon Molecular mass i'll write as Mr \[Unknown~mole~of~C =\frac{ mole ~of~Methane \times Mr~of ~ C }{ Mr~of~Methane }\] and so you get mole of C in Methane do same thing but now using mole of Ethane and get the answer

OpenStudy (somy):

Step 3 now you have mole of C in methane and in ethane so add both moles up and get total mole and then \[No.~of~C~atoms= total~mole~of~C \times (6.02 \times 10^{23}) \]

OpenStudy (somy):

6.02 * 10 ^23 is Avogadro constant btw

OpenStudy (somy):

is it clear i hope?

OpenStudy (anonymous):

thanks @Somy

OpenStudy (somy):

no problem :)